OpenStudy (anonymous):

which expression is equivalent to c-0.25c

OpenStudy (shamim):

1-0.25=?

OpenStudy (asad786):

75

OpenStudy (asad786):

what do you mean by c-25

OpenStudy (asad786):

what is c your trying to ask right

OpenStudy (anonymous):

0.75c

OpenStudy (asad786):

@RaniaZaki

OpenStudy (welshfella):

these are like terms so c - 0.25c = 0.75c

OpenStudy (asad786):

exactly :)

OpenStudy (asad786):

whas there any answer choices

OpenStudy (anonymous):

x>5x-8

OpenStudy (anonymous):

-4x>-8 -4x/-4<-8/-4 x<2 A=]-infinity ; 2[